# arXiv:2603.02085v1[hep-ph] 02 Mar 2026

### v1

@@ Abstract
The relation between the effective degrees of freedom (EDOF) and the trace anomaly is studied in the hadron resonance gas (HRG) model. If we regard the thermodynamical relation as the evolution equation and define the EDOF as $P/T^{4}$, where $P$ and $T$ are the pressure and the temperature, respectively, we obtain the equation which relates to the trace anomaly. The structure of the equation resembles that of the so called c-theorem in the two dimensional conformal field theory which asserts that the EDOF should not increase as the energy scale parameter decreases. There is a stationary point where the trace anomaly (modified trace anomaly) vanishes, and the scale symmetry is restored. To investigate the limiting temperature of the HRG model with the excluded volume effects, we consider two types of the c-theorem like conditions for the EDOF. The first condition requires that the EDOF should not decrease when $T$ increases. This condition is equivalent to the condition that the trace anomaly (modified trace anomaly) should not be negative. The second condition requires that the EDOF should be convex downwards as a function of $T$. It is found that the first condition gives the limiting temperature of the HRG model with the excluded volume effect which is much higher than the crossover transition temperature obtained by the lattice QCD calculation and, at zero baryon number density, is close to the transition temperature in the pure gluonic theory, while the second one gives the limiting temperature which almost coincides with the one obtained by using the normalized baryon number fluctuation in the previous work and is consistent with the critical point predicted by the lattice QCD calculation.
@@ I Introduction
The deconfinement of quarks is a mysterious phenomenon whose mechanism is not yet clear. Due to the strong interaction, at low temperature and low density, quarks are confined in hadrons, namely baryons and mesons. The lattice QCD (LQCD) calculations indicate that, at high temperature, quarks are deconfined and form quark matter. However, at $\mu=0$ where $\mu$ is the baryon number chemical potential, the LQCD calculations predict that the transition from hadron matter to quark matter is not a discontinuous phase transition but a continuous crossover transition Aoki et al. (2006). It is not well understood how quarks break free from hadrons during the crossover transition. Several new intermediate phases between the hadron phase and the ideal quark gluon plasma (QGP) phase are proposed; e.g., see Refs. Pisarski (2006); Glozman (2015); Cohen and Glozman (2024); Fujimoto et al. (2025) and references therein.
@@
At low temperature, the equation of state obtained by LQCD calculations can be well reproduced by the hadron resonance gas (HRG) model. In the HRG model, the ideal gas approximation is usually used. In such an approximation, as temperature increases, the effective degrees of freedom (EDOF) increase rapidly, and the HRG results deviate from the LQCD ones. The ideal gas model has no mechanism that limits the model itself, since there is no interaction in the model, although the resonances themselves are produced by the interaction. To obtain the limiting temperature of hadrons, in particular, the repulsive interaction that makes hadrons unstable is needed. The repulsive interaction reduces the EDOF of hadron matter and makes hadron matter thermodynamically unfavorable. The more the EDOF of freedom of hadron increases, the stronger the repulsive effects between hadrons become.
@@
In Ref. Oshima et al. (2026), the baryon number fluctuation was used to investigate the limiting temperature of the baryon gas model. Figure 1 shows $T$-dependence of the normalized baryon number fluctuation $\chi_{2}^{\rm B}/T^{2}$ at $\mu=0$ in the HRG models with EVE and without EVE; for details of the models, see Sec. III. In the HRG model without EVE, $\chi_{2}^{\rm B}/T^{2}$ increases monotonically as $T$ increases. In the HRG model with EVE, it reaches its maximum at $T_{\chi,\rm max}=0.195$ GeV. We see that the result in the HRG model with EVE is consistent with that in LQCD up to $T_{\chi,\rm max}$. This indicates that $T_{\chi,\rm max}$ is the limiting temperature of the baryon gas model. In Ref. Oshima et al. (2026), the $\mu$-dependence of $T_{\chi,\rm max}$ was also analyzed. It was found that the LQCD predicted critical point (CP) is located almost on the curve of $T_{\rm max}(\mu)$ in the $\mu$-$T$ plane. However, the mechanism behind this coincidence is not yet clear.
@@
The dimensionless quantity $\chi_{2}^{\rm B}/T^{2}$ is related to the EDOF. In two dimensional conformal field theory, using the renormalization group method and the entropic theory, it was shown that the EDOF does not increase as the energy scale of the system decreases Zamolodchikov (1986); Casini and Huerta (2007). It is called the c-theorem. In the c-theorem, the c-function $c(r)$, which represents the EDOF of the theory and converges to the central charge at the conformal fixed point, is defined as a function of the length scale parameter $r$ of the system, and $\frac{\partial c(r)}{\partial r}\leq 0$ is shown. This inequality ensures that the EDOF does not increase as $r$ increases. Inversely, it may be natural that the effective degrees of freedom do not decrease as the energy scale such as temperature increases. In this paper, to investigate the limiting temperature of the HRG model, we require this condition for the normalized pressure which is defined as EDOF in our picture.
@@
We also consider a stronger condition. Figure 2 shows the pressure $P$ divided by $T^{4}$, namely, the EDOF $g_{\mu}(T)$ in our picture, obtained by the LQCD calculation when $\mu=0$. The EDOF has an inflection point as a function of $T$ when $T\sim 0.175$ GeV. At the inflection point, $T^{5}$-normalized trace anomaly has its maximum. The EDOF is convex downwards above the inflection point and is convex upwards below it. In our simple picture, the region below the inflection point is regarded as the hadron phase, and the region above the point is regarded as the high temperature phase such as quark gluon plasma (QGP) or the exotic new phase; see Fig. 3. Hence, the EDOF of the hadron phase is convex downwards. We require this condition for the EDOF in the HRG model.
@@ II Effective degrees of freedom and the trace anomaly
${\partial g_{T}(\mu)\over{\partial\mu}}={\epsilon-3P-Ts\over{\mu^{5}}}={\Delta-Ts\over{\mu^{5}}}$ (6)
@@ IV Effective degrees of freedom and trace anomaly in Hadron resonance gas model
Fig. 5 shows the $T^{5}$-normalized trace anomaly $\Delta/T^{5}$ as a function of $T$ in the HRG model with and without EVE, when $\mu=0$. The baryon contribution $\Delta_{\rm B}/T^{5}$ increases monotonically as it increases in the HRG model without EVE, while it has its maximum at $T=0.197$ GeV in the HRG model with EVE. It should be also noted that $\Delta/T^{5}$ has its local maximum at $T=0.041$ GeV in the HRG model. This local maximum is induced by the pion contribution. Fig. 6 shows the pion contribution to $\Delta/T^{5}$ in the HRG model. It is clear that this local peak is induced by the pion contribution. Since the pion is the pseudo Nambu-Goldstone boson and its mass is much smaller than that of the other hadrons, the pion contribution $\Delta_{\pi}/T^{5}$ has local peak in low temperature. This property is not unnatural, hence, we do not regard that this peak structure indicates the limitation of the HRG model. This peak structure remains in the model with EVE to mesons, since it locates in quite low temperature region. For the EVE to mesons, see Sec. V.
Fig. 7: The $T$-dependence of the $\beta^{3}\Delta$ ($\Delta/T^{3}$) at $\mu=0$. The dashed and dash-dot-dotted lines show the results in the HRG model with EVE and without EVE, respectively. The solid and dash-dotted lines show the baryon contribution $\beta^{3}\Delta_{\rm M}$ in the HRG model with and without EVE, respectively. The dotted line shows result in the ideal massless three flavor QGP with (without) the gluon contributions. The squares with error bar show the LQCD results in Ref. Borsanyi et al. (2012a).
@@
Figure 10 and  11 show $g_{\gamma}(\beta)$ and $\Delta/T^{3}$ as a function of $T$, respectively, when $\gamma=-6$. Since the baryon contribution is dominant at $\gamma=-\mu/T=-6$, the dashed-dotted and the dash-dot-dotted lines almost coincide in the HRG model without EVE. In the HRG model with EVE, the difference between the solid and dashed lines is not small, since the baryon contribution is suppressed by the excluded volume effects. In the case of the HRG model without EVE, $g_{\gamma}(\beta)$ increases monotonically as $T$ increases, while it has its local maximum (minimum) at $T=0.169$ ($T=0.210$ GeV) where $\Delta$ vanishes, and its baryon contribution $g_{\gamma,\rm B}(\beta)$ has its maximum at $T=0.157$ GeV where $\Delta_{\rm B}$ vanishes in the HRG model with EVE, and the maximum point of $\beta^{3}\Delta_{\rm B}$, $T=1/\beta=0.120$ GeV, corresponds to the inflection point of $g_{\gamma,\rm B}(\beta)$. Since the baryon contribution is a main contribution at large $\gamma$, $\beta^{3}\Delta$ also have their maxima at $T=0.121$ GeV.
Fig. 12: The dash-dot-dotted line shows the $T_{\rm zero}$ where $\Delta_{\rm B}-\mu n_{\rm B}$ vanishes in the HRG model with EVE. The solid (dash-dotted) line shows the temperature $T_{\rm max}$ ($T_{\chi\rm max}$) where $(\Delta_{\rm B}-\mu n_{\rm B})/T^{5}$ ($\chi_{2}^{\rm B}/T^{2}$) has its maximum for the fixed value of $\mu$. The solid squares with the error bar show the LQCD crossover line taken from Ref.  Borsanyi et al. (2020). The open squares with the error bar show the temperature $T_{\rm max,L}$ where $\Delta/T^{5}$ has its maximum obtained by using the LQCD data in Ref.  Borsanyi et al. (2012a). The crosses with error bar shows the LQCD predicted critical point taken from Ref. Shah et al. (2024).
@@
In Fig. 12, the curves of $T_{\rm max}(\mu)$ and $T_{\chi,\rm max}(\mu)$ where $(\Delta_{\rm B}-\mu n_{\rm B})/T^{5}$ and $\chi_{2}^{\rm B}/T^{2}$ have their maximum for the fixed value of $\mu$ in the HRG model with EVE, respectively, are also shown. The $T_{\rm max}$ is the limiting temperature of the HRG model with EVE obtained by the strong condition (13). We see that the critical point predicted by LQCD is located almost on these two curves. The two curves almost coincide up to CP and then deviate. In the large $\mu$ region, the curve of $T_{\rm max}$ lies below that of $T_{\chi,\rm max}$. This fact supports our assumption that, at least up to CP, the $T_{\rm max}(\mu)$ curve represents the limiting temperature of the baryon gas model.
@@
We also show the temperature $T_{\rm max,L}^{\prime\prime}$ where the $T^{3}$-normalized trace anomaly $\Delta/T^{3}(=\beta^{3}\Delta)$ obtained by using the average value of $\Delta/T^{4}$ of the LQCD data in Ref. Borsanyi et al. (2012a). Since the lattice data is discretized in $T$, we regard the temperature interval as the systematic error bar. The $T_{\rm max,L}^{\prime}$ is located just above $T_{\rm max}^{\prime}$ in the HRG model with EVE. Note that $T_{\rm max,L}^{\prime\prime}$ is obtained for the fixed value of $\mu$, not for $\gamma$. Hence, we also show $T_{\rm max}^{\prime\prime}$ where $\beta^{3}\Delta$ has its maximum for the fixed value of $\mu$ in the HRG model with EVE. The $T_{\rm max,L}^{\prime}$ is also somewhat larger than $T_{\rm max}^{\prime\prime}$.
@@ V Excluded volume effects in the meson gas
$M_{i}(T)$ $=$ ${g_{{\rm M},i}\over{2\pi^{2}}}\int_{0}^{\infty}dpp^{2}e^{-\sqrt{p^{2}+m_{{\rm M},i}^{2}}/T},$ (40)
@@
Fig. 17: The dash-dot-dotted (upper dashed) line shows the $T_{\rm zero2}$ ($T_{\rm zeroM}$) where $\Delta-\mu n_{\rm B}$ ($\Delta_{M}$) vanishes in the HRG model with EVE2. The dotted (lower dashed) line shows the temperature $T_{\rm max2}$ ($T_{\rm maxM}$) where $(\Delta-\mu n_{\rm B})/T^{5}$ ($\Delta_{\rm M}/T^{5}$) has its maximum for the fixed value of $\mu$. The thin solid line in the low $T$ and the small $\mu$ region is the local maximum or minimum of $(\Delta-\mu n_{\rm B})/T^{5}$. The meaning of the solid line and the symbols are the same as in Fig. 12.
Figure 17 shows the $\mu$-dependence of the temperature $T_{\rm zero2}$ where $(\Delta-\mu n_{\rm B})$ vanishes in the HRG model with EVE2. The $T_{\rm zero2}$ is the limiting temperature of the HRG model with EVE obtained by the c-theorem like condition (10). At small $\mu$, the curve of $T_{\rm zero2}$ is much higher than the LQCD crossover line and, at $\mu=0$, is close to the transition temperature $T_{\rm d}\sim 0.285$ GeV in the pure gluonic theory. Borsanyi et al. (2022); Giusti et al. (2025); Fujimoto et al. (2025). At $\mu=\mu_{\rm CP}$, $T_{\rm zero2}$ is much higher than $T_{\rm CP}$. The temperature $T_{\rm zerozM}$ ($T_{\rm maxM}$) where $\Delta_{\rm M}$ vanishes (has its maximum) is constant and is 0.301 (0.199) GeV, since $\Delta_{\rm M}$ does not depend on $\mu$.
@@
$(\mu_{\rm CRP},T_{\rm CRP})=(0.594~{\rm GeV},0.115~{\rm GeV}).$ (41)
This indicates that the limiting temperature of the HRG model with EVE contains some information about the critical point, although the model has no mechanism for the chiral phase transition.

### v2

# Effective degrees of freedom, trace anomaly and c-theorem like condition in the hadron resonance gas model

@@ Abstract
The relation between the effective degrees of freedom (EDOF) and the trace anomaly is studied in the hadron resonance gas (HRG) model. If we regard the thermodynamical relation as the evolution equation and define the EDOF as $P/T^{4}$, where $P$ and $T$ are the pressure and the temperature, respectively, we obtain the equation which relates to the trace anomaly. The structure of the equation resembles that of the so-called c-theorem, which asserts that the EDOF should not increase as the energy scale parameter decreases, in the two dimensional conformal field theory. There is a stationary point where the trace anomaly (modified trace anomaly) vanishes, and the scale symmetry is restored. To investigate the limiting temperature of the HRG model with the excluded volume effects, we consider two types of the c-theorem like conditions for the EDOF. The first condition requires that the EDOF should not decrease when $T$ increases. This condition is equivalent to the condition that the trace anomaly (modified trace anomaly) should not be negative. The second condition requires that the EDOF should be convex downwards as a function of $T$. It is found that the first condition gives the limiting temperature of the HRG model with the excluded volume effect which is much higher than the crossover transition temperature obtained by the lattice QCD calculation and, at zero baryon number density, is close to the transition temperature in the pure gluonic theory, while the second one gives the limiting temperature which almost coincides with the one obtained by using the normalized baryon number fluctuation in the previous study and is consistent with the critical point predicted by the lattice QCD calculation.
@@ I Introduction
The deconfinement of quarks is a mysterious phenomenon whose mechanism is not clear yet. Due to the strong interaction, at low temperature and low density, quarks are confined in hadrons, namely baryons and mesons. The lattice QCD (LQCD) calculations indicate that, at high temperature, quarks are deconfined and form quark matter. However, at $\mu=0$ where $\mu$ is the baryon number chemical potential, the LQCD calculations predict that the transition from hadron matter to quark matter is not a discontinuous phase transition but a continuous crossover transition Aoki et al. (2006). It is not well understood how quarks break free from hadrons during the crossover transition. Several new intermediate phases between the hadron phase and the ideal quark gluon plasma (QGP) phase are proposed; e.g., see Refs. Pisarski (2006); Glozman (2015); Cohen and Glozman (2024); Fujimoto et al. (2025) and references therein.
@@
At low temperature, the equation of state obtained by LQCD calculations can be well reproduced by the hadron resonance gas (HRG) model. In the HRG model, the ideal gas approximation is usually used. In such an approximation, as temperature increases, the effective degrees of freedom (EDOF) increase rapidly, and the HRG results deviate from the LQCD ones. The ideal gas model has no mechanism that limits the model itself, since there is no interaction in the model, although the resonances themselves are produced by the interaction. To obtain the limiting temperature of hadrons, in particular, the repulsive interaction that makes hadrons unstable is needed. The repulsive interaction reduces the EDOF of hadron matter and makes hadron matter thermodynamically unfavorable. The more the effective degrees of freedom of hadron increase, the stronger the repulsive effects between hadrons become.
@@
In Ref. Oshima et al. (2026), the baryon number fluctuation was used to investigate the limiting temperature of the baryon gas model. Figure 1 shows $T$-dependence of the normalized baryon number fluctuation $\chi_{2}^{\rm B}/T^{2}$ at $\mu=0$ in the HRG models with EVE and without EVE; for details of the models, see Sec. III. In the HRG model without EVE, $\chi_{2}^{\rm B}/T^{2}$ increases monotonically as $T$ increases. In the HRG model with EVE, it reaches its maximum at $T_{\chi,\rm max}=0.195$ GeV. We see that the result in the HRG model with EVE is consistent with that in LQCD up to $T_{\chi,\rm max}$. This indicates that $T_{\chi,\rm max}$ is the limiting temperature of the baryon gas model. In Ref. Oshima et al. (2026), the $\mu$-dependence of $T_{\chi,\rm max}$ was also analyzed. It was found that the LQCD predicted critical point (CP) is located almost on the curve of $T_{\rm max}(\mu)$ in the $\mu$-$T$ plane. This limiting temperature is also related to the $R=0$ criterion, where $R$ is the scalar curvature in the thermodynamic geometry Castorina et al. (2018, 2019, 2020a); Zhang et al. (2020); Castorina et al. (2020b); Murgana et al. (2024). However, the mechanism behind this coincidence is not clear yet.
@@
The dimensionless quantity $\chi_{2}^{\rm B}/T^{2}$ is related to the EDOF. In two dimensional conformal field theory, using the renormalization group method and the entropic theory, it was shown that the EDOF does not increase as the energy scale of the system decreases Zamolodchikov (1986); Casini and Huerta (2007). It is called the c-theorem. In the c-theorem, the c-function $c(r)$, which represents the EDOF of the theory and converges to the central charge at the conformal fixed point, is defined as a function of the length scale parameter $r$ of the system, and $\frac{\partial c(r)}{\partial r}\leq 0$ is shown. This inequality ensures that the EDOF does not increase as $r$ increases. Inversely, it may be natural that the effective degrees of freedom do not decrease as the energy scale such as temperature increases. In this paper, to investigate the limiting temperature of the HRG model, we assume this condition for the normalized pressure which is defined as EDOF in our picture.
@@
We also consider a stronger condition. Figure 2 shows the pressure $P$ divided by $T^{4}$, namely, the EDOF $g_{\mu}(T)$ in our picture, obtained by the LQCD calculation when $\mu=0$. The EDOF has an inflection point as a function of $T$ when $T\sim 0.175$ GeV. At the inflection point, $T^{5}$-normalized trace anomaly has its maximum. The EDOF is convex downwards above the inflection point and is convex upwards below it. In our simple picture, the region below the inflection point is regarded as the hadron phase, and the region above the point is regarded as the high temperature phase such as quark gluon plasma (QGP) or the exotic new phase; see Fig. 3. (To say more precisely, since the transition is crossover at $\mu=0$, the low (high) temperature phase is composed mainly of hadron (QGP/exotic new matter). ) Hence, the EDOF of the hadron phase is convex downwards. We assume this condition for the EDOF in the HRG model.
@@ II Effective degrees of freedom and the trace anomaly
${\partial g_{T}(\mu)\over{\partial\mu}}={\epsilon-3P-Ts\over{\mu^{5}}}={\Delta-Ts\over{\mu^{5}}}.$ (6)
@@ IV Effective degrees of freedom and trace anomaly in Hadron resonance gas model
Fig. 5 shows the $T^{5}$-normalized trace anomaly $\Delta/T^{5}$ as a function of $T$ in the HRG model with and without EVE, when $\mu=0$. The baryon contribution $\Delta_{\rm B}/T^{5}$ increases monotonically as it increases in the HRG model without EVE, while it has its maximum at $T=0.197$ GeV in the HRG model with EVE. It should be also noted that $\Delta/T^{5}$ has its local maximum at $T=0.041$ GeV in the HRG model. This local maximum is induced by the pion contribution. Fig. 6 shows the pion contribution to $\Delta/T^{5}$ in the HRG model. It is clear that this local peak is induced by the pion contribution. Since the pion is the pseudo Nambu-Goldstone boson and its mass is much smaller than those of the other hadrons, the pion contribution $\Delta_{\pi}/T^{5}$ has local peak in low temperature. This property is not unnatural, hence, we do not regard that this peak structure indicates the limitation of the HRG model. This peak structure remains in the model with EVE to mesons, since it locates in quite low temperature region. For the EVE to mesons, see Sec. V.
Fig. 7: The $T$-dependence of the $\beta^{3}\Delta$ ($\Delta/T^{3}$) at $\mu=0$. The dashed and dash-dot-dotted lines show the results in the HRG model with EVE and without EVE, respectively. The solid and dash-dotted lines show the baryon contribution $\beta^{3}\Delta_{\rm B}$ in the HRG model with and without EVE, respectively. The dotted line shows result in the ideal massless three flavor QGP with (without) the gluon contributions. The squares with error bar show the LQCD results in Ref. Borsanyi et al. (2012a).
@@
Figure 10 and  11 show $g_{\gamma}(\beta)$ and $\Delta/T^{3}$ as a function of $T$, respectively, when $\gamma=-6$. Since the baryon contribution is dominant at $\gamma=-\mu/T=-6$, the dashed-dotted and the dash-dot-dotted lines almost coincide in the HRG model without EVE. In the HRG model with EVE, the difference between the solid and dashed lines is not small, since the baryon contribution is suppressed by the excluded volume effects. In the case of the HRG model without EVE, $g_{\gamma}(\beta)$ increases monotonically as $T$ increases, while it has its local maximum (minimum) at $T=0.169$ ($T=0.210$ GeV) where $\Delta$ vanishes, and its baryon contribution $g_{\gamma,\rm B}(\beta)$ has its maximum at $T=0.157$ GeV where $\Delta_{\rm B}$ vanishes in the HRG model with EVE, and the maximum point of $\beta^{3}\Delta_{\rm B}$, $T=1/\beta=0.120$ GeV, corresponds to the inflection point of $g_{\gamma,\rm B}(\beta)$. Since the baryon contribution is a main contribution at large $\gamma$, $\beta^{3}\Delta$ also have their maxima at $T\sim 0.120$ GeV.
Fig. 12: The dash-dot-dotted line shows the $T_{\rm zero}$ where $\Delta_{\rm B}-\mu n_{\rm B}$ vanishes in the HRG model with EVE. The solid (dash-dotted) line shows the temperature $T_{\rm max}$ ($T_{\chi,\rm max}$) where $(\Delta_{\rm B}-\mu n_{\rm B})/T^{5}$ ($\chi_{2}^{\rm B}/T^{2}$) has its maximum for the fixed value of $\mu$. The solid squares with the error bar show the LQCD crossover line taken from Ref.  Borsanyi et al. (2020). The open squares with the error bar show the temperature $T_{\rm max,L}$ where $\Delta/T^{5}$ has its maximum obtained by using the LQCD data in Ref.  Borsanyi et al. (2012a). The crosses with error bar shows the LQCD predicted critical point taken from Ref. Shah et al. (2024).
@@
In Fig. 12, the curves of $T_{\rm max}(\mu)$ and $T_{\chi,\rm max}(\mu)$ where $(\Delta_{\rm B}-\mu n_{\rm B})/T^{5}$ and $\chi_{2}^{\rm B}/T^{2}$ have their maximum for the fixed value of $\mu$ in the HRG model with EVE, respectively, are also shown. The $T_{\rm max}$ is the limiting temperature of the HRG model with EVE obtained by the strong condition (13). We see that the critical point predicted by the LQCD is located almost on these two curves. The two curves almost coincide up to CP and then deviate. In the large $\mu$ region, the curve of $T_{\rm max}$ lies below that of $T_{\chi,\rm max}$. This fact supports our assumption that, at least up to CP, the $T_{\rm max}(\mu)$ curve represents the limiting temperature of the baryon gas model.
@@
We also show the temperature $T_{\rm max,L}^{\prime\prime}$ where the $T^{3}$-normalized trace anomaly $\Delta/T^{3}(=\beta^{3}\Delta)$ obtained by using the average value of $\Delta/T^{4}$ of the LQCD data in Ref. Borsanyi et al. (2012a). Since the lattice data is discretized in $T$, we regard the temperature interval as the systematic error bar. The $T_{\rm max,L}^{\prime}$ is located just above $T_{\rm max}^{\prime}$ in the HRG model with EVE. Note that $T_{\rm max,L}^{\prime\prime}$ is obtained for the fixed value of $\mu$, not for $\gamma$. Hence, we also show $T_{\rm max}^{\prime\prime}$ where $\beta^{3}\Delta$ has its maximum for the fixed value of $\mu$ in the HRG model with EVE. The $T_{\rm max,L}^{\prime\prime}$ is also somewhat larger than $T_{\rm max}^{\prime\prime}$.
@@ V Excluded volume effects in the meson gas
$M_{i}(T)$ $=$ ${g_{{\rm M},i}\over{2\pi^{2}}}\int_{0}^{\infty}dpp^{2}e^{-\sqrt{p^{2}+m_{{\rm M},i}^{2}}/T}.$ (40)
@@
Fig. 17: The dash-dot-dotted (upper dashed) line shows the $T_{\rm zero2}$ ($T_{\rm zero2M}$) where $\Delta-\mu n_{\rm B}$ ($\Delta_{M}$) vanishes in the HRG model with EVE2. The dotted (lower dashed) line shows the temperature $T_{\rm max2}$ ($T_{\rm max2M}$) where $(\Delta-\mu n_{\rm B})/T^{5}$ ($\Delta_{\rm M}/T^{5}$) has its maximum for the fixed value of $\mu$. The thin solid line in the low $T$ and the small $\mu$ region is the local maximum or minimum of $(\Delta-\mu n_{\rm B})/T^{5}$. The meaning of the solid and dash-dotted lines and the symbols are the same as in Fig. 12.
Figure 17 shows the $\mu$-dependence of the temperature $T_{\rm zero2}$ where $(\Delta-\mu n_{\rm B})$ vanishes in the HRG model with EVE2. The $T_{\rm zero2}$ is the limiting temperature of the HRG model with EVE obtained by the c-theorem like condition (10). At small $\mu$, the curve of $T_{\rm zero2}$ is much higher than the LQCD crossover line and, at $\mu=0$, is close to the transition temperature $T_{\rm d}\sim 0.285$ GeV in the pure gluonic theory. Borsanyi et al. (2022); Giusti et al. (2025); Fujimoto et al. (2025). At $\mu=\mu_{\rm CP}$, $T_{\rm zero2}$ is much higher than $T_{\rm CP}$. The temperature $T_{\rm zero2M}$ ($T_{\rm max2M}$) where $\Delta_{\rm M}$ vanishes (has its maximum) is constant and is 0.301 (0.199) GeV, since $\Delta_{\rm M}$ does not depend on $\mu$.
@@
$(\mu_{\rm CRP},T_{\rm CRP})=(0.594~{\rm GeV},0.115~{\rm GeV}),$ (41)
and is very close to the CP. This indicates that the limiting temperature of the HRG model with EVE contains some information about the critical point, although the model has no mechanism for the chiral phase transition.
@@
However, it should be remarked that this local peak structure induced by pion disappears if we use a dimensionless scale parameter $t=\log{(T/T_{0})}$ where $T_{0}$ is a constant parameter with the same dimension as $T$. Such a dimensionless scale parameter is usually used in the renormalization group equation. In this case, the evolution equation is given by
${\partial g_{\mu}(T)\over{\partial t}}=T{\partial g_{\mu}(T)\over{\partial T}}$ $=$ ${\Delta-\mu n_{\rm B}\over{T^{4}}}$ (42)
Note that this modification does not change $T_{\rm zero2}$ where the right hand side of the equation vanishes. However, the temperature where the right hand side has its (local) maximum changes or disappears. Figure 18 is the same as Fig. 17 but the evolution equation (42) is used. We see that the local peak structure induced by pion disappears. On the other hand, qualitative characteristics of the the global maximum structure is retained although $T_{\rm max2t}$ is somewhat larger than $T_{\rm max2}$ in the small $\mu$ region. It should be also remarked that $T_{\rm max2t}$ is close to $T_{\rm max2,Lt}$ obtained by using the LQCD result Borsanyi et al. (2012a). In this meaning, the $(t,\mu)$ coordinate system may be more preferable than $(T,\mu)$.
Fig. 18: The dash-dot-dotted (upper dashed) line shows the $T_{\rm zero2}$ ($T_{\rm zeroM}$) where $\Delta-\mu n_{\rm B}$ ($\Delta_{M}$) vanishes in the HRG model with EVE2. The dotted (solid, lower dashed) line shows the temperature $T_{\rm max2t}$ ($T_{\rm maxt}$$T_{\rm max2Mt}$) where $(\Delta-\mu n_{\rm B})/T^{4}$ $($$\Delta_{\rm B}-\mu n_{\rm B})/T^{4}$$\Delta_{\rm M}/T^{4}$) has its maximum for the fixed value of $\mu$. The open squares with the error bar show the temperature $T_{\rm max,Lt}$ where $\Delta/T^{4}$ has its maximum obtained by using the LQCD data in Ref.  Borsanyi et al. (2012a). The meaning of the dash-dotted line and the other symbols are the same as in Fig. 12.